# VeriCoder: Enhancing LLM-Based RTL Code Generation through Functional Correctness Validation

### v1

@@ Abstract
Recent advances in Large Language Models (LLMs) have sparked growing interest in applying them to Electronic Design Automation (EDA) tasks, particularly Register Transfer Level (RTL) code generation. While several RTL datasets have been introduced, most focus on syntactic validity rather than functional validation with tests, leading to training examples that compile but may not implement the intended behavior. We present VeriCoder, a model for RTL code generation fine-tuned on a dataset validated for functional correctness. This fine-tuning dataset is constructed using a novel methodology that combines unit test generation with feedback-directed refinement. Given a natural language specification and an initial RTL design, we prompt a teacher model (GPT-4o-mini) to generate unit tests and iteratively revise the RTL design based on its simulation results using the generated tests. If necessary, the teacher model also updates the tests to ensure they comply with the natural language specification. As a result of this process, every example in our dataset is functionally validated, consisting of a natural language description, an RTL implementation, and passing tests. Fine-tuned on this dataset of over 125,000 examples, VeriCoder achieves state‑of‑the‑art metrics in functional correctness on VerilogEval and RTLLM, with relative gains of up to 71.7% and 27.4% respectively. An ablation study further shows that models trained on our functionally validated dataset outperform those trained on functionally non-validated datasets, underscoring the importance of high-quality datasets in RTL code generation.
@@
RTL code generation, large language model.
@@ I Introduction
In this work, we introduce VeriCoder, a model for RTL code generation fine-tuned on a high-quality dataset consisting of 125,777 examples that has been validated for functional correctness. To construct this dataset, we develop a novel pipeline that combines unit test generation with feedback-directed refinement guided by a teacher LLM (GPT-4o-mini). Given a natural language specification and an initial RTL implementation, the teacher model first generates a unit test. If the RTL code fails the simulation, the model iteratively revises the design based on the observed error messages. When needed, the unit test is also updated to better reflect the intended functionality described by the specification. This process continues until the design passes simulation or a retry limit is reached. The resulting fine-tuning dataset consists of over 125k validated triples: a natural language specification, a correct RTL design, and a self-checking unit test.
@@
  • We introduce VeriCoder [^1] We plan to open source the dataset, model, and code upon acceptance., an RTL code generation model fine‑tuned on a dataset validated for functional correctness. On the VerilogEval and RTLLM benchmarks, VeriCoder achieves state‑of‑the‑art performance among open‑source fine‑tuned models, yielding relative pass@k gains of up to 71.7% and 27.4% over the prior best.
  • We develop a dataset augmentation pipeline that combines unit test generation with feedback-directed refinement guided by a teacher LLM. This yields, to the best of our knowledge, the largest fine-tuning dataset to date with functional validation, consisting of over 125k validated triples of natural language specifications, RTL designs, and passing tests.
@@ II-A Language Modeling and Fine-Tuning
II-B RTL Code Datasets
@@
Mining open-source RTL designs is a common strategy for dataset construction. VeriGen [26] compiles Verilog modules from GitHub and textbooks into a structured corpus using automated syntax checks. BetterV [24] collects Verilog modules from the internet and then filters designs based on coding style and syntactic validity. Other works [8, 29, 30] adopt similar methodologies for sourcing and preprocessing RTL code.
@@
A recent work, OpenLLM-RTL [31], explores the idea of using LLMs to generate assertions, producing a functionally verified dataset of 7k examples. While sharing the same goal of improving functional correctness in fine-tuning datasets, our work takes a different approach by generating unit tests for validation. Our final dataset contains over 125k examples, by far the largest functionally validated RTL dataset.
@@ III-C Algorithm and Prompts
[…]
@@
While we cannot guarantee that every design in the augmented dataset is functionally correct under all possible inputs, the inclusion of unit tests makes it substantially more robust than prior approaches that rely solely on syntactic checking. We consider this a practical and scalable step toward constructing higher-quality fine-tuning datasets for RTL generation.
@@ IV-C Benchmarks and Metrics
[…]
@@ V-A Main Evaluation Results
VeriCoder achieves state-of-the-art results across two RTL code generation benchmarks, outperforming all previously released open-source fine-tuned models. On VerilogEval-Machine, VeriCoder attains a pass@1 accuracy of 55.7%, representing a 19.8 percentage point improvement over the best prior model, OriGen. On VerilogEval-Human, it reaches 38.3%, exceeding OriGen by 16.0 percentage points. Across all evaluated $k$-shot settings ($k{=}1,5,10$), VeriCoder consistently maintains its lead on the Human split. On the RTLLM benchmark, VeriCoder achieves 79.3% syntax correctness and 48.3% functional correctness, surpassing OriGen’s 51.7% and 37.9%, respectively. In conclusion, VeriCoder delivers relative improvements of up to 71.7% on VerilogEval and 27.4% on RTLLM in pass@k accuracy, surpassing the previous state-of-the-art model on both benchmarks.
@@
Despite the observed improvements, a substantial performance gap persists between VeriCoder and the strongest large models. For instance, o3-mini attains 66.4% on VerilogEval Pass@1 compared to VeriCoder’s 55.7%. DeepSeek-R1 achieves 69.1% on human-graded Pass@5, versus VeriCoder’s 49.2%. Commercial LLMs such as GPT-4o reach perfect 100.0% Syntax-VCS validity and 69.0% functional correctness, while VeriCoder records 79.3% and 48.3%, respectively. Despite the performance gap, open-source lightweight models offer compelling advantages. They provide transparency, allow for local deployment, and ensure intellectual property protection, i.e., capabilities that are particularly important for RTL design workflows where security, customizability, and integration into existing toolchains are critical.
@@ V-B Ablation Study of Dataset
Across all metrics, we observe a consistent improvement as dataset quality increases. On the VerilogEval benchmark, the base model achieves 46.8% Pass@5. Fine-tuning on the unvalidated dataset raises performance to 53.5%, while our validated dataset further improves it to 55.8%. For RTLLM syntax correctness, the trend is similar: 69.0% for the base model, 75.9% for the unvalidated version, and 79.3% when trained on validated data. Functional correctness sees even more significant improvement, rising from 41.4% (base) to 44.8% (unvalidated) and ultimately to 48.3% (validated).
@@ VII Conclusion
Recent advances in Large Language Models (LLMs) have opened new possibilities for Electronic Design Automation (EDA), particularly in RTL code generation. However, most existing datasets emphasize syntactic validity while overlooking functional correctness, which limits the effectiveness of fine-tuned models. We introduce VeriCoder, a model fine-tuned on a dataset with 125,000 examples that is validated for functional correctness. This dataset is constructed using a feedback-directed refinement pipeline guided by a teacher LLM, which generates and iteratively updates both RTL designs and unit tests until the design passes simulation. The resulting dataset consists of functionally validated triples comprising a natural language specification, an RTL implementation, and a passing test. Fine-tuned on this dataset, VeriCoder achieves state-of-the-art results on two established RTL benchmarks, yielding relative improvements of up to 71.7% on VerilogEval and 27.4% on RTLLM. An ablation study confirms the impact of functional validation on model performance, underscoring the importance of high-quality training data. Looking ahead, future work may incorporate formal verification and reinforcement learning to further improve models’ performance in AI-assisted hardware design.

### v2

@@ Abstract
Recent advances in Large Language Models (LLMs) have sparked growing interest in applying them to Electronic Design Automation (EDA) tasks, particularly Register Transfer Level (RTL) code generation. While several RTL datasets have been introduced, most focus on syntactic validity rather than functional validation with tests, leading to training examples that compile but may not implement the intended behavior. We present VeriCoder, a model for RTL code generation fine-tuned on a dataset validated for functional correctness. This fine-tuning dataset is constructed using a novel methodology that combines unit test generation with feedback-directed refinement. Given a natural language specification and an initial RTL design, we prompt a teacher model (GPT-4o-mini) to generate unit tests and iteratively revise the RTL design based on its simulation results using the generated tests. If necessary, the teacher model also updates the tests to ensure they comply with the natural language specification. As a result of this process, every example in our dataset is functionally validated, consisting of a natural language description, an RTL implementation, and passing tests. Fine-tuned on this dataset of 125,777 examples, VeriCoder achieves state‑of‑the‑art metrics in functional correctness on VerilogEval and RTLLM, with relative gains of up to 71.7% and 27.4%, respectively. An ablation study further shows that models trained on our functionally validated dataset outperform those trained on functionally non-validated datasets, underscoring the importance of high-quality datasets in RTL code generation. Our code, data, and models are publicly available at https://github.com/Anjiang-Wei/VeriCoder
@@
RTL, Code Generation, Large Language Model.
@@ I Introduction
In this work, we introduce VeriCoder, a model for RTL code generation fine-tuned on a high-quality dataset consisting of 125,777 examples that has been validated for functional correctness [^1] While functional correctness is not fully guaranteed, we manually reviewed 100 randomly sampled examples and found that 92% of the generated RTL code correctly matches the corresponding natural language descriptions.. To construct this dataset, we develop a novel pipeline that combines unit test generation with feedback-directed refinement guided by a teacher LLM (GPT-4o-mini). Given a natural language specification and an initial RTL implementation, the teacher model first generates a unit test. If the RTL code fails the simulation, the model iteratively revises the design based on the observed error messages. When needed, the unit test is also updated to better reflect the intended functionality described by the specification. This process continues until the design passes simulation or a retry limit is reached. The resulting fine-tuning dataset consists of 125,777 validated triples: a natural language specification, a correct RTL design, and a self-checking unit test.
@@
  • We introduce VeriCoder, an RTL code generation model fine‑tuned on a dataset validated for functional correctness. On the VerilogEval and RTLLM benchmarks, VeriCoder achieves state‑of‑the‑art performance among open‑source fine‑tuned models, yielding relative pass@k gains of up to 71.7% and 27.4% over the prior best.
  • We develop a dataset augmentation pipeline that combines unit test generation with feedback-directed refinement guided by a teacher LLM. This yields, to the best of our knowledge, the largest fine-tuning dataset to date with functional validation, consisting of 125,777 validated triples of natural language specifications, RTL designs, and passing tests.
@@ II-A Language Modeling and Fine-Tuning
II-B Related Work on RTL Code Generation
@@
Mining open-source RTL designs is a common strategy for dataset construction. VeriGen [26] compiles Verilog modules from GitHub and textbooks into a structured corpus using automated syntax checks. BetterV [24] collects Verilog modules from the internet and then filters designs based on coding style and syntactic validity. CraftRTL [29] augments fine-tuning data with non-textual code representations, injecting synthetic errors derived from intermediate model checkpoints into open-source Verilog code. Other works [8, 30, 31] adopt similar methodologies for sourcing and preprocessing RTL code.
@@
A recent work, OpenLLM-RTL [32], explores the idea of using LLMs to generate assertions, producing a functionally verified dataset of 7k examples. While sharing the same goal of improving functional correctness in fine-tuning datasets, our work takes a different approach by generating unit tests for validation. Our final dataset contains over 125,777 examples—the largest functionally validated RTL dataset to date.
Beyond data collection and synthesis techniques, several works explore other methods to enhance RTL code generation quality. ScaleRTL [33] emphasizes reasoning by generating intermediate traces and leveraging test-time compute through iterative self-reflection. DeepRTL [34] adopts curriculum learning guided by multi-level natural language summaries. VeriSeek [35] applies reinforcement learning with feedback derived from AST-level similarity between LLM outputs and reference designs. AutoVCoder [36] incorporates retrieval-augmented generation (RAG), dynamically supplying relevant Verilog snippets to the model. CodeV [37] extends generation capabilities to tasks such as fill-in-the-middle (FIM). Our work adopts standard supervised fine-tuning while focusing on constructing a large-scale, functionally validated dataset. Our approach is complementary and orthogonal to existing techniques.
@@ III-C Algorithm and Prompts
4:Define: $\mathit{GenTestTpl}\leftarrow$ prompt template for test generation
5:  $\mathit{RefineTpl}\leftarrow$ prompt template for iterative refinement
@@
While functional correctness under all possible inputs cannot be guaranteed, the inclusion of unit tests makes our augmented dataset substantially more robust than prior approaches that rely solely on syntactic checking. We view this as a practical and scalable step toward building higher-quality fine-tuning datasets for RTL generation. To assess quality, we manually reviewed 100 randomly sampled examples and found that 92% of the generated RTL code correctly matched the corresponding natural language descriptions.
@@ IV-C Benchmarks and Metrics
$\text{Pass@}k=\mathbb{E}\left[1-\frac{\binom{n-c}{k}}{\binom{n}{k}}\right$]
@@ V-A Main Evaluation Results
VeriCoder achieves state-of-the-art results across two RTL code generation benchmarks, outperforming all previously released open-source fine-tuned models. On VerilogEval-Machine, VeriCoder attains a pass@1 accuracy of 55.7%, representing a 19.8 percentage point improvement over the best prior model, OriGen. On VerilogEval-Human, it reaches 38.3%, exceeding OriGen by 16.0 percentage points. Across all evaluated $k$-shot settings (k=1,5,10k{=}1,5,10), VeriCoder consistently maintains its lead on the Human split. On the RTLLM benchmark, VeriCoder achieves 79.3% syntax correctness and 48.3% functional correctness, surpassing OriGen’s 51.7% and 37.9%, respectively. In conclusion, VeriCoder delivers relative improvements of up to 71.7% on VerilogEval and 27.4% on RTLLM in pass@k accuracy, surpassing the previous state-of-the-art model on both benchmarks.
@@
Despite the observed improvements, a substantial performance gap persists between VeriCoder and the strongest large models. For instance, o3-mini attains 66.4% on VerilogEval Pass@1 compared to VeriCoder’s 55.7%. DeepSeek-R1 achieves 69.1% on human-graded Pass@5, versus VeriCoder’s 49.2%. Commercial LLMs such as GPT-4o reach a perfect 100.0% Syntax-VCS validity and 69.0% functional correctness, while VeriCoder records 79.3% and 48.3%, respectively. Despite the performance gap, open-source lightweight models offer compelling advantages. They provide transparency, allow for local deployment, and ensure intellectual property protection, i.e., capabilities that are particularly important for RTL design workflows where security, customizability, and integration into existing toolchains are critical.
@@ V-B Ablation Study of Dataset
Across all metrics, we observe a consistent improvement as dataset quality increases. On the VerilogEval benchmark (covering both Machine and Human subsets), the base model achieves 46.8% Pass@5. Fine-tuning on the unvalidated dataset raises performance to 53.5%, while our validated dataset further improves it to 55.8%. For RTLLM syntax correctness, the trend is similar: 69.0% for the base model, 75.9% for the unvalidated version, and 79.3% when trained on validated data. Functional correctness sees even more significant improvement, rising from 41.4% (base) to 44.8% (unvalidated) and ultimately to 48.3% (validated).
@@ VII Conclusion
Recent advances in Large Language Models (LLMs) have opened new possibilities for Electronic Design Automation (EDA), particularly in RTL code generation. However, most existing datasets emphasize syntactic validity while overlooking functional correctness, which limits the effectiveness of fine-tuned models. We introduce VeriCoder, a model fine-tuned on a dataset with 125,000 examples that is validated for functional correctness. This dataset is constructed using a feedback-directed refinement pipeline guided by a teacher LLM, which generates and iteratively updates both RTL designs and unit tests until the design passes simulation. The resulting dataset consists of functionally validated triples comprising a natural language specification, an RTL implementation, and a passing test. Fine-tuned on this dataset, VeriCoder achieves state-of-the-art results on two established RTL benchmarks, yielding relative improvements of up to 71.7% on VerilogEval and 27.4% on RTLLM. An ablation study confirms the impact of functional validation on model performance, underscoring the importance of high-quality training data. Future work may explore formal verification and reinforcement learning to further advance AI-assisted hardware design.
Acknowledgment
We thank Samantha Archer, Yao Hsiao, Mohammad Rahmani Fadiheh and Subhasish Mitra for their discussions. This work was partially supported by a Google Research Award.